Again, some fog in the west in the morning, truck temperature +3F, seemed to be linked to the Deer Creek reservoir.  Left a lot of fluffy frost on the DCS tower...

Spent the day at DCS with Clayton, Sebastian, and later Matt and Dexter.  Removed everything except the primary fence, the AC service, and one argumentative fence post and ground rod, preventing the trailer from being pulled.  Also some hardware (e.g. 2 darkhorses and the Rohn sections) is still there.

Extracted all soils with a few nicks, but presumed working.  The heat flux plate was off of horizontal by perhaps 5 degrees and the TP01 foil was a bit wrinkled.

At the very end of the day, ran out to CC to soak the lock with Kroil.  It now works with the combination we expected.